# Exploring the Mechanisms Behind Non Aromatic Fluorescence with Density Functional Tight Binding Method

### 20 Jan 2024 v2

@@ Abstract
Recent experimental findings reveal non-conventional fluorescence emission in systems devoid of conjugated bonds or aromatic compounds, termed Non-Aromatic Fluorescence (NAF). This phenomenon is exclusive to aggregated or solid states, remaining absent in monomeric solutions. Previous studies on small system models in vacuum emphasize that the carbonyl stretching mode, along with the strong interaction of Short Hydrogen Bonds (SHBs), remains the primary vibrational mode explaining NAF in these systems. This work investigates NAF in the crystal structure of L-pyroglutamine-ammonium, comparing it with the related non-fluorescent amino acid L-glutamine. Employing Density Functional Tight Binding (DFTB) theory in combination with Non-Adiabatic Molecular Dynamics (NAMD) and the Mixed Quantum/Molecular Mechanics (QM/MM) approach, we simulate larger model systems and consider the effects of the surrounding environment. Our results on more realistic systems align with previous reports using higher-level theories. Moreover, our study demonstrates the efficacy of DFTB in elucidating the intricacies inherent to NAF, underscoring the qualitative insight into this novel optical property. Beyond contributing to the ongoing exploration of NAF, this work forms the basis for delving further into the complexities characterizing non-aromatic systems within the fluorescence paradigm.
@@ 1 Introduction
Over the last decade, a growing body of experimental evidence has revealed the emergence of non-conventional fluorescence signals in biomolecular systems. In contrast to our standard spectroscopic intuition attributing fluorescence in biological systems to aromatic or conjugated groups, recent intriguing experiments have shown that non-aromatic systems exhibit inherent absorption and fluorescence in the UV-Visible range1, 2, 3, 4. These fluorescent signals appear to be enhanced in aggregated forms, analogous to the aggregation-induced emission observed in organic materials5, 6 Noteworthy examples include amyloid structures7, 8, 9, individual amino acids10, 11, compounds derived from amino acids12, 10, poly-amidoamines13 and small organic molecules14, 15.
@@
Over the last few years, our group has been leading theoretical and computational efforts to investigate the spectroscopic origins of this phenomenon9, 16, 12, 18, 1. We have recently investigated NAF in model amyloid proteins9, 18 and L-glutamine/L-glutamine-derived16, 12 amino-acids combining both experiments and theory. Employing Trajectory Surface Hopping (TSH)19, 20, simulations at a Time Dependent Density Functional Theory (TD-DFT) level21, we have shown that carbonyl groups (CO) appear to play a critical role in modulating the observation of NAF in a wide variety of systems.
@@
Upon an initial photo-excitation to the first excited state $S_{1}$, which is mainly localized around the CO group, strong inter-molecular interactions involving hydrogen bonds within certain biological aggregates hinder the stretching of the CO bond. This leads to a restriction of the non-radiative relaxation pathways, subsequently boosting the fluorescence quantum yield. The CO stretching mode is, at the same time, strongly coupled to other vibrational modes involving the amide bond which appear to be modulated by local environmental effects. Our previous research has been based on dimer model systems carved out of crystal structures; in these studies we have accounted for the steric effect of the surrounding molecules by incorporating specific constraints in our simulations12, 9, 16. The computational demands associated with simulating the photophysics of these systems make TSH calculations extremely intensive. Furthermore, the inherent variability in the electronic relaxation pathways can introduce slow convergence problems. Consequently, a substantial number of trajectories becomes essential to derive meaningful and reliable results. We have recently employed hybrid Quantum-Classical Mechanical (QM/MM) approaches to enable a realistic description of the environment effects on the fluorescence.16 However, the computationally tractable sizes of the quantum mechanical region are severely limited. These limitations have sparked our interest in exploring alternative methods to effectively tackle these challenges. In this regard, Density Functional Tight-Binding (DFTB)22, 23, an approximate method based on Density Functional Theory (DFT), has emerged as a promising approach. DFTB achieves computational efficiency through a combination of the two-center approximation, pre-tabulated integrals, and the use of a minimal basis set24, 25, 26. This accelerates DFTB calculations by a factor of 2-3 compared to traditional DFT25. Moreover, the pre-tabulation of integrals at DFT level ensures the precision of DFTB in predicting electronic properties for molecular, nano, and periodic systems, effectively emulating DFT27, 28, 29, 30, 31, 32. This framework has also been extended to include the modeling of excited state dynamics leading to the Time-Dependent Density Functional Tight-Binding method (TD-DFTB). These extensions include both the frequency33, 34(TD-DFTB) and the time domain35, 36 (real-time TD-DFTB). In addition, TD-DFTB has been coupled to the Ehrenfest approach35, 36 and TSH37, 38, 39 to perform non-adiabatic dynamics.
@@
The goal of this work is to comprehensively assess the potential of the TD-DFTB method in investigating the photophysics underlying NAF phenomenon employing the DFTB+ code40. The primary focus of our investigation is on amino-acid crystals which have recently been studied in our group using both ab-initio electronic methods (TD-DFT) and experiments. It was found that L-glutamine, characterized by regular hydrogen bond interactions, is a non-fluorescent system. However, following a chemical transformation into L-pyro-ammonium (pyroglutamic acid complexed with an ammonium ion), the system undergoes a transition to a short hydrogen bond configuration and displays a remarkable fluorescent signal. (Panels a and b of Figure 1 show the molecular structure of these systems.). By comparing and contrasting the mechanisms obtained with TD-DFTB to our previous work, we demonstrate that the DFTB approach correctly captures the non-radiative decay pathways occurring in L-glutamine as well as the origins of fluorescence in L-pyro-ammonium. With this validation in hand, we use TD-DFTB to understand the role of environmental effects using a QM/MM approach allowing for an accurate characterization of the experimentally observed Stoke-Shift between the excitation and emission spectra in this system. We also highlight aspects of non-radiative decay mechanisms which are different from our previous studies paving the way for future directions and developments needed to use Tight-Binding approaches to study non-aromatic fluorescence.
@@ 2 Theoretical Approaches
The objective of this section is to provide a background on the various theories and techniques employed in our methodologies to study non-aromatic fluorescence. In this section, we will summarize some of the most important aspects of DFTB, as well as its Time-Dependent extension to the frequency domain (TD-DFTB), with a special focus on the implementations made in the DFTB+ code40. Additionally, we will describe the key aspects of non-adiabatic dynamics in the context of Trajectory Surface Hopping (TSH), with a specific emphasis on the different analytical expressions used to calculate the non-radiative probability decay.
@@ 2.1 Density Functional Tight-Binding (DFTB and TD-DFTB)
$E^{DFT}[\rho^{0}+\delta\rho]\approx E^{0}[\rho^{0}]+E^{1\text{st}}[\rho^{0},\delta\rho]+E^{2\text{nd}}[\rho^{0},(\delta\rho)^{2}]+E^{3\text{rd}}[\rho^{0},(\delta\rho)^{3}].$ (1)
@@
$E^{DFTB1}=E^{0}+E^{1\text{st}}=\sum_{A>B}E_{AB}^{rep}+\sum^{occ.}_{i}n_{i}\langle\psi_{i}|\mathcal{H}[\rho^{0}]|\psi_{i}\rangle,$ (2)
where $\langle\psi_{i}|$ is the $i^{\text{th}}$ molecular orbital in a linear combination of atomic orbitals (LCAO), $n_{i}$ is the occupation and $\mathcal{H}[\rho^{0}]$ is the Hamiltonian operator in a two-center approximation.
In the second-order DFTB (DFTB2) method, the $E^{2nd}$ term is incorporated through a monopole approximation23, where the electron density fluctuations are approximated as a sum of atomic Slater-type spherical charge densities:
$E^{DFTB2}=E^{DFTB1}+\frac{1}{2}\sum_{AB}\gamma_{AB}\Delta q_{A}\Delta q_{B}.$ (3)
Here, $\Delta q_{A}$ is the Mulliken charge on atom A and $\gamma_{AB}$ represents the electron interaction of two Slater-type spherical charge densities on atoms A and B. The inclusion of the second term requires a self-consistent solution since the Mulliken charges depend on the molecular orbitals. This provides a more accurate description of charge-transfer and charge-charge interactions.
@@
$E^{DFTB3}=E^{DFTB2}+\frac{1}{3}\sum_{AB}\Gamma_{AB}\Delta q_{A}^{2}\Delta q_{B},$ (4)
@@
Within the DFTB+40 code there are several available approaches to determine excited states. We employed the Casida formalism21 which within the framework of DFTB, follows the same ansatz as in the original version of TD-DFT34, 33. The electronic excitation energy $\omega_{I}$ in TD-DFTB can be obtained by solving the following eigenvalue problem:
@@
Moreover, the incorporation of long-range corrections into the two-electron integrals have also been implemented in the DFTB+ code46 with the aim to provide a reduction in the self-interaction error in DFT (and also present in DFTB) yielding a correct behaviour of the asymptotic potential. This implementation along with a new reparametrization of the Slater-Koster parameters enables a more accurate description of the excited states, mainly for those which involves charge transfer excitations.47
@@ 2.2 Trajectory Surface Hopping (TSH)
Trajectory Surface Hopping (TSH) is one of the most widely employed techniques to model non-adiabatic dynamics20. The non-radiative relaxation is a non-equilibrium process and therefore the dynamics strongly depends on the initial conditions. For this reason, the simulation of a large number of thermally distributed independent trajectories, with varying initial conditions becomes essential48, 49. In each TSH trajectory, the electronic states are propagated using quantum mechanics, while the nuclear motion is handled classically employing the forces coming from a single potential energy surface from the quantum method. For more information on the theoretical and practical details, the reader is referred to previous studies20, 50. Bellow, we highlight specific details relevant to our present work.
@@
The $g_{ij}$ terms are the non adiabatic coupling (NAC) elements, one of the key components in non-adiabatic dynamics. Several alternatives approaches have been developed in order to avoid the explicit determination of the non adiabatic coupling elements51, 52, 53, since sometimes there are no analytical expressions available or because they are computationally expensive to determine for large systems. One of the most popular in this context is the Landau-Zener (LZSH) approximation54:
$P_{ij}(t)=exp\left(-\dfrac{\pi}{2\hbar}\sqrt{\dfrac{\Delta E_{ij}^{3}(t)}{\dfrac{d^{2}}{dt^{2}}\Delta E_{ij}(t)}}\right)$ (8)
where $\Delta E_{ij}$ is the adiabatic energy gap55, 50. Within the context of our TD-DFTB simulations, we employ the LZSH approximation instead of Tully’s method due to its easy implementation into our computational framework.
@@ 3 Computational Details
In this section, we summarize the computational protocols applied to L-glutamine (L-gln) and L-pyroglutamine ammonium (L-pyro-amm) to simulate the ground-state molecular dynamics in periodic boundary conditions, the determination of the absorption spectra, the excited states surface hopping dynamics and finally the QM/MM simulations with DFTB/TD-DFTB. In the present study, we employed the DFTB+ package40 for the electronic structure calculations. The Genesis56, 57 software was used to perform classical and QM/MM simulations.
@@ 3.1 Ground State DFTB Molecular Dynamics
Ab-initio molecular dynamics simulations of both L-gln and L-pyro-amm were conducted using DFTB in the ground-state in periodic boundaries conditions (PBC) using the gamma point. The initial coordinates for both systems were obtained from previously resolved crystallographic structures (CSD 116931658 and 198155112). We employed third order corrections in the Hamiltonian with the corresponding Slater Koster parameters 3ob59 as well as D3 Grimme corrections42. Molecular dynamics simulations were conducted for a total of 600 ps for each system using a timestep of 0.5 fs and within the NVT ensemble using the Berendsen thermostat with a coupling of 0.1 $ps^{-1}$. For all the analysis, the first 20 ps were discarded as equilibration. As will be shown later, we compare some of the ground-state structural properties obtained from DFTB to those obtained with DFT. Details of these simulations can be found in the Computational Methods Section of a previous work in our group12.
@@ 3.2 Absorption Spectra
These calculations were carried out within the TD-DFTB method. To account for the correct description of the excited states, we also incorporated Long Range Corrections (LC) with the appropriate Slater-Koster files47. A total of 100 states were calculated for each snapshot employing the Casida algorithm. The final absorption spectrum for each system was generated by averaging the frames, and each individual band was broadened using a Gaussian function with a Full Width at Half Maximum (FWHM) of 5 nm.
3.3 Excited State Molecular Dynamics in Gas Phase
@@
Our protocol to generate the initial conditions consists of performing 6 ns long classical molecular dynamics simulations using the Generalized Amber Force Field (GAFF)60 for both model systems with a timestep of 1 fs and a thermostat constant of 1 $ps^{-1}$. From these long simulations, a total of 600 configurations were selected with a spacing of 10 ps. For each of these 600 configurations, ground-state simulations were continued using DFTB as described earlier in section 3.1, for 2 ps each at 300K.
TSH simulations were started with the system in the electronic state $S_{1}$ corresponding to the initial photo-absorption. The dynamics were then performed in the NVE ensemble for 250 fs using a timestep of 0.5 fs with TD-DFTB as described in section 3.2. At each timestep, the probability of non-radiative decay was calculated using the LZSH algorithm (Equation 8).
As noted earlier in the Introduction, one of the goals of this work is to compare and contrast the underlying non-radiative decay mechanisms associated with non-aromatic fluorescence obtained with TD-DFT in our earlier work16 to TD-DFTB. It is therefore important to recall two important differences. Firstly, the use of TD-DFTB expands the statistics of TSH trajectories by a factor of 3 (600 vs 200) in the case of L-gln and a factor of 6 for L-pyro-amm (600 vs 100). Secondly, as already indicated earlier, the TD-TBDFT hopping events between different states are performed using LZSH (Equation 8) while in our previous TD-DFT simulations, were conducted with FSSH (Equation 7). Despite this methodological difference, the comparison between the two approaches remains valid, as both theories have been extensively tested and compared in several works61, 62, 63, demonstrating consistent results.

### 09 Apr 2024 v3

# Exploring the Mechanisms Behind Non Aromatic Fluorescence with the Density Functional Tight Binding Method

@@ Abstract
Recent experimental findings reveal non-conventional fluorescence emission in biological systems devoid of conjugated bonds or aromatic compounds, termed Non-Aromatic Fluorescence (NAF). This phenomenon is exclusive to aggregated or solid states, remaining absent in monomeric solutions. Previous studies focused on small model systems in vacuum show that the carbonyl stretching mode along with strong interaction of short hydrogen bonds (SHBs) remain the primary vibrational mode explaining NAF in these systems. In order to simulate larger model systems taking into account the effects of the surrounding environment, in this work we propose using the density functional tight-binding (DFTB) method in combination with non-adiabatic molecular dynamics (NAMD) and the mixed quantum/molecular mechanics (QM/MM) approach. We investigate the mechanism behind NAF in the crystal structure of L-pyroglutamine-ammonium, comparing it with the related non-fluorescent amino acid L-glutamine. Our results extend our previous findings to more realistic systems, demonstrating the efficiency and robustness of the proposed DFTB method in the context of NAMD in biological systems. Furthemore, due to its inherent low computational cost, this method allows for a better sampling on the non-radiative events at the conical intersection which is crucial for a complete understanding of this phenomenon. Beyond contributing to the ongoing exploration of NAF, this work paves the way for future application of this method in more complex biological systems such as amyloid aggregates, biomaterials and non-aromatic proteins.
@@ 1 Introduction
Over the last decade, a growing body of experimental evidence has revealed the emergence of non-conventional fluorescence signals in biomolecular systems. In contrast to conventional spectroscopic intuition attributing fluorescence in biological systems to aromatic or conjugated groups, recent experiments have shown that non-aromatic systems exhibit inherent absorption and fluorescence in the UV-visible range1, 2, 3, 4. These fluorescent signals appear to be enhanced in aggregated forms, analogous to the aggregation-induced emission observed in organic materials5, 6. Noteworthy examples include amyloid structures7, 8, 9, individual amino acids10, 11, compounds derived from amino acids12, 10, poly-amidoamines13 and small organic molecules14, 15.
@@
Over the last few years, our group has been leading theoretical and computational efforts to investigate the spectroscopic origins of this phenomenon9, 16, 12, 18, 1. We have recently investigated NAF in model amyloid proteins9, 18 and L-glutamine/L-glutamine-derived16, 12 amino-acids combining both experiments and theory. Employing trajectory surface hopping (TSH)19, 20 simulations in combination with time-dependent density functional theory (TD-DFT) level21, we have shown that carbonyl groups (CO) appear to play a critical role in modulating the observation of NAF in a wide variety of systems. Upon an initial photo-excitation to the first excited state $S_{1}$, which is mainly localized around the CO group, strong inter-molecular interactions involving hydrogen bonds within certain biological aggregates hinder the stretching of the CO bond. This leads to a restriction of the non-radiative relaxation pathways, thus boosting the fluorescence quantum yield. The CO stretching mode is strongly coupled to the proton transfer and deplanarization modes in the amide group, which are also restrained due to environment effects16.
@@
Our previous research has been based on dimer model systems carved out of crystal structures; in these studies we have accounted for the steric effect of the surrounding molecules by incorporating specific constraints in our simulations12, 9, 16. The computational demands associated with simulating the photophysics of these systems make TSH calculations extremely challenging. Furthermore, the inherent variability in the electronic relaxation pathways can introduce problems with convergence. Consequently, the inclusion of a substantial number of trajectories becomes essential to derive meaningful and reliable results. We have recently employed hybrid Quantum-Classical Mechanical (QM/MM) approaches to enable a realistic description of environmental effects on the fluorescence.16 However, these approaches are computationally limited by the size of the relevant quantum-mechanical subsystem. These limitations have sparked our interest in exploring alternative methods to effectively tackle these challenges. In this regard, the density functional tight-binding (DFTB) method22, 23, a semi-empirical approach based on density functional theory (DFT), has emerged as a promising avenue. DFTB achieves computational efficiency through a combination of the two-center approximation, pre-tabulated integrals, and the use of a minimal basis set24, 25, 26. This accelerates DFTB calculations by a factor of 2-3 order of magnitude compared to traditional DFT25. Moreover, the pre-tabulation of integrals at DFT level ensures the precision of DFTB in predicting electronic properties for molecular, nano, and periodic systems, effectively emulating DFT27, 28, 29, 30, 31, 32. This framework has also been extended to include the modeling of excited state dynamics leading to the time-dependent density functional tight-binding method (TD-DFTB). These extensions include both the frequency33, 34(TD-DFTB) and the time domain35, 36 (real-time TD-DFTB). In addition, TD-DFTB has been coupled to the Ehrenfest approach35, 36 and TSH37, 38, 39 to perform non-adiabatic dynamics with semiclassical ion motion.
@@
The goal of the present work is to comprehensively assess the potential of the TD-DFTB method in investigating the photophysics underlying NAF phenomenon employing the DFTB+ code40. The primary focus of our investigation is on amino-acid crystals which have recently been studied in our group using both ab-initio electronic methods (TD-DFT) and experiments. It was found that L-glutamine, characterized by regular hydrogen bond interactions, is a non-fluorescent system. However, following a chemical transformation into L-pyroglutamine-ammonium (L-pyro-ammonium, pyroglutamic acid complexed with an ammonium ion), the system undergoes a transition to a short hydrogen bond configuration and displays a remarkable fluorescent signal. Panels a and b of Figure 1 show the molecular structure of these systems. By comparing and contrasting the mechanisms obtained with TD-DFTB to our previous work, we demonstrate that the DFTB approach correctly captures the non-radiative decay pathways occurring in L-glutamine as well as the origins of fluorescence in L-pyro-ammonium. With this validation in hand, we use TD-DFTB to understand the role of environmental effects using a QM/MM approach allowing for an accurate characterization of the experimentally observed Stokes shift between the excitation and emission spectra in this system. We also highlight aspects of non-radiative decay mechanisms which are different from our previous studies paving the way for future directions and developments needed to use Tight-Binding approaches to study non-aromatic fluorescence.
@@ 2 Theoretical Approaches
The objective of this section is to provide a background on the various theories and techniques employed in our methodologies to study non-aromatic fluorescence. In this section, we will summarize some of the most important aspects of DFTB, as well as its Time-Dependent extension to the frequency domain (TD-DFTB), with a special focus on the implementations in the DFTB+ code40. Additionally, we will describe the key aspects of non-adiabatic dynamics in the context of Trajectory Surface Hopping (TSH), with a specific emphasis on the different analytical expressions used to calculate the non-radiative probability decay.
@@ 2.1 Density Functional Tight-Binding (DFTB and TD-DFTB)
$E^{\text{DFT}}[\rho^{0}+\delta\rho]\approx E^{0}[\rho^{0}]+E^{1\text{st}}[\rho^{0},\delta\rho]+E^{2\text{nd}}[\rho^{0},(\delta\rho)^{2}]+E^{3\text{rd}}[\rho^{0},(\delta\rho)^{3}].$ (1)
@@
$E^{\text{DFTB1}}=E^{0}+E^{1\text{st}}=\sum_{A>B}E_{AB}^{\text{rep}}+\sum^{\text{occ.}}_{i}n_{i}\langle{\psi_{i}}|\mathcal{H}[\rho^{0}]|\psi_{i}\rangle,$ (2)
where $|\psi_{i}\rangle$ is the $i^{\text{th}}$ molecular orbital in a linear combination of atomic orbitals (LCAO), $n_{i}$ is the occupation and $\mathcal{H}[\rho^{0}]$ is the Hamiltonian operator in a two-center approximation.
In the second-order DFTB (DFTB2) method, the $E^{2\text{nd}}$ term is incorporated through a monopole approximation23, where the electron density fluctuations are approximated as a sum of atomic Slater-type spherical charge densities:
$E^{\text{DFTB2}}=E^{\text{DFTB1}}+\frac{1}{2}\sum_{AB}\gamma_{AB}\Delta q_{A}\Delta q_{B}.$ (3)
Here, $\Delta q_{A}$ is the Mulliken charge on atom $A$ and $\gamma_{AB}$ represents the electron interaction of two Slater-type spherical charge densities on atoms $A$ and $B$. The inclusion of the second-order term requires a self-consistent solution since the Mulliken charges depend on the molecular orbitals. This provides a more accurate description of charge-transfer and charge-charge interactions.
@@
$E^{\text{DFTB3}}=E^{\text{DFTB2}}+\frac{1}{3}\sum_{AB}\Gamma_{AB}\Delta q_{A}^{2}\Delta q_{B},$ (4)
@@
Within the DFTB+40 code there are several available approaches to determine excited states. We employed the Casida formalism21 within the framework of DFTB, following the same ansatz as in the original version of TD-DFT34, 33. The electronic excitation energy $\omega_{I}$ in TD-DFTB can be obtained by solving the following eigenvalue problem:
@@
Moreover, the incorporation of Long-range Corrections (LC) into the two-electron integrals have also been implemented in the DFTB+ code46 with the aim to provide a reduction in the self-interaction error in DFT (and also present in DFTB) yielding a correct behaviour of the asymptotic potential. This implementation along with a new reparametrization of the Slater-Koster parameters enables a more accurate description of the excited states, mainly for those which involves charge transfer excitations47, 48. It is worth noting that in previous literature, the term TD-LC-DFTB denotes the incorporation of LC into TD equations. However, in our current work, we will use the term TD-DFTB since all calculations within this theory were conducted with LC.
@@ 2.2 Trajectory Surface Hopping (TSH)
Trajectory Surface Hopping (TSH) is one of the most widely employed techniques to model non-adiabatic dynamics20. The non-radiative relaxation is a non-equilibrium process and therefore the dynamics strongly depends on the initial conditions. For this reason, the simulation of a large number of thermally distributed independent trajectories, with varying initial conditions becomes essential49, 50. In each TSH trajectory, the electronic states are propagated using quantum mechanics, while the nuclear motion is handled classically employing the forces coming from a single potential energy surface from the quantum method. For more information on the theoretical and practical details, the reader is referred to previous studies20, 51. Bellow, we highlight specific details relevant to our present work.
@@
where $C_{i}$ are the electronic coefficients of the potential energy surfaces $i$ and the $g_{ij}$ terms are the non adiabatic coupling (NAC) elements, one of the key components in non-adiabatic dynamics. Several alternatives approaches have been developed in order to avoid the explicit determination of the non adiabatic coupling elements52, 53, 54, since sometimes there are no analytical expressions available or because they are computationally expensive to determine for large systems. One of the most popular in this context is the Landau-Zener (LZSH) approximation55:
$P_{ij}(t)=\exp{\left(-\dfrac{\pi}{2\hbar}\sqrt{\dfrac{\Delta E_{ij}^{3}(t)}{\dfrac{d^{2}}{dt^{2}}\Delta E_{ij}(t)}}\right)},$ (8)
where $\Delta E_{ij}$ is the adiabatic energy gap56, 51. Within the context of our TD-DFTB simulations, we employ the LZSH approximation instead of Tully’s method due to its easy implementation into our computational framework.
@@ 3 Computational Details
In this section, we summarize the computational protocols applied to L-glutamine (L-gln) and L-pyroglutamine ammonium (L-pyro-amm) to simulate the ground-state molecular dynamics in periodic boundary conditions, the determination of the absorption spectra, the excited states surface hopping dynamics and finally the QM/MM simulations with DFTB and TD-DFTB. In the present study, we employed the DFTB+ package40 for the electronic structure calculations. The Genesis57, 58 software was used to perform classical and QM/MM simulations.
@@ 3.1 Ground State DFTB Molecular Dynamics
[…]
@@ 3.2 Absorption Spectra
These calculations were carried out within the TD-DFTB method without PBC, which corresponds to the simulation of a cluster of the unit cell for each system surrounded by vacuum. To account for the correct description of the excited states, we also incorporated Long-range Corrections (LC) with the appropriate Slater-Koster files47, 48. A total of 100 states were calculated for each snapshot employing the Casida algorithm. The final absorption spectrum for each system was generated by averaging the frames, and each individual transition was broadened using a Gaussian function with a Full Width at Half Maximum (FWHM) of 5 nm.
3.3 Non Adiabatic Molecular Dynamics in Gas Phase
@@
[…]
As noted earlier in the Introduction, one of the goals of this work is to compare and contrast the underlying non-radiative decay mechanisms associated with non-aromatic fluorescence obtained with TD-DFTB to the one reported in our earlier work16, employing TD-DFT with the functional PBE0 and 6-31G as the basis set.
It is therefore important to recall two important differences: (i) the use of TD-DFTB expands the statistics of TSH trajectories by a factor of 3 (600 vs 200) in the case of L-gln and a factor of 6 for L-pyro-amm (600 vs 100), (ii) as already indicated earlier, the TD-DFTB hopping events between different states are performed using LZSH (Equation 8) while in our previous TD-DFT simulations, were conducted with FSSH (Equation 7). Despite this methodological difference, the comparison between the two approaches remains valid, as both theories have been extensively tested and compared in several works62, 63, 64, demonstrating consistent results.